# Learn With Imagination: Safe Set Guided State-wise Constrained Policy Optimization

### v3

@@ D.1 Preliminaries
Proof.
Multiply both sides of (36) by $(I-\gamma P_{\pi})$ and take the inner product with the vector $f\in\mathbb{R}^{|\hat{\mathcal{S}}|}$. ∎
Combining Lemma 7 with (35), we obtain the following, for any function $f$ and any policy $\pi$:
@@
After bringing (44), $\left\|\dagger{\delta}_{f}^{\pi^{\prime}}\right\|_{\infty}$ into (43), then substract $\underset{\begin{subarray}{c}\hat{s}\sim\dot{d}^{\pi}\\ a\sim\pi\\ \hat{s}^{\prime}\sim P\end{subarray}}{E}[\delta_{f}(\hat{s},a,\hat{s}^{\prime})]$, the bounds are obtained. The lower bound leads to (40), and the upper bound leads to (41). ∎
Lemma 9.
@@
Proof.
Begin with the bounds from lemma 8 and bound the divergence $D_{TV}(\dot{d}^{\pi^{\prime}}||\dot{d}^{\pi})$ by lemma 9. ∎
@@ D.2 Safety invariance in expectation between S-3PO policies
Corollary 3 (SCPO Update Constraint Satisfaction).
@@
Proof.
The choice of $f=\hat{V}_{D}^{\pi},g=D$ in lemma 10 leads to following inequality:
@@
where $\mathcal{\hat{J}}_{D}(\pi)=\mathbb{E}_{\tau\thicksim\pi}\Bigg{[}\sum_{t=0}^{H}\gamma^{t}D\big{(}\hat{s}_{t},a_{t},\hat{s}_{t+1}\big{)}\Bigg{]}$, need to distinguish from $\mathcal{J}_{D}(\pi)$. And $\hat{V}_{D}^{\pi},\hat{A}_{D}^{\pi}$ are also the discounted version of $V_{D}^{\pi}$ and $A_{D}^{\pi}$. Note that according to Lemma 10 one can only get this the inequality holds when $\gamma\in(0,1)$.
@@
where $\bar{d}^{\pi}=\sum_{t=0}^{H}P(\hat{s}_{t}=\hat{s}|\pi)$. Thus, following the inequality (33), the Corollary 3 is proofed.
D.3 Proof of Theorem 2
If $s_{t}\in\mathcal{\bar{S}}\in\mathcal{S}_{S}^{D}$, then $\mathcal{J}_{D}(\pi_{t})\leq 0$ and $\phi(s_{t})\leq 0$. According to Corollary 3, we know that if policy $\pi_{t}$ is updated by solving (33), then we have $\mathcal{J}_{D}(\pi_{t+1})\leq 0$ which means:
@@ D.3 Proof of Theorem 2
Appendix E Proof of Theorem 3
@@ E.2 Trust Region Update Performance
Proof.
The choice of $f=V_{\pi},g=R$ in lemma 10 leads to following inequality:
@@
At this point, the lemma 12 is proved.
E.3 Proof of Theorem 3
Proof.
If ISSA will constantly be triggered, the safeguard should be treated as a component of the environment, which indicates the environment is non-stationary for policy $\pi$ after safeguard is disabled. According to Algorithm 1, we assume that ISSA will not be triggered after $k_{safe}$ updates, meaning the environment is stationary for policy $\pi$ with or without safeguard. Therefore, we can infer S-3PO worst performance degradation after $k_{safe}$ updates following the trust region results of finite-horizon MDPs. Utilizing Lemma 12 and the relationship between the total variation divergence and the KL divergence, we have:
@@ E.3 Proof of Theorem 3
Due to Lemma 11, if two policies are related with Equation 57, they are related with the following optimization:
@@ F.1 Environment Settings
To scale our method to high dimensional link robots. We additionally adopt Walker shown in 17 as a bipedal robot ($\mathcal{A}\subseteq\mathbb{R}^{10}$) in our experiments.
@@ F.2 Policy Settings
F.3 Metrics Comparison
In Tables 5(c), 6(c) and 7, we report all the $9$ results of our test suites by three metrics:
@@ F.3 Metrics Comparison
Both the evaluation performance and training performance are reported based on the above metrics. Besides, we also report the ISSA triger times as ISSA performance of TRPO-ISSA and S-3PO. All of the metrics were obtained from the final epoch after convergence. Each metric was averaged over two random seed. The evaluation performance curves of all experiments are shown in Figures 19, 25 and 31, the training performance curves of all experiments are shown in Figures 21, 27 and 33 and the ISSA performance curves of all experiments are shown in Figures 23, 29 and 35
@@
(a) Point_1Hazard
(b) Point_4Hazard
(c) Point_8Hazard
Figure 19: Evaluation performance of Point_Hazard
(a) Point_1Hazard
(b) Point_4Hazard
(c) Point_8Hazard
Figure 21: Training performance of Point_Hazard
(a) Point_1Hazard
(b) Point_4Hazard
(c) Point_8Hazard
Figure 23: ISSA performance of Point_Hazard
(a) Point_1Pillar
(b) Point_4Pillar
(c) Point_8Pillar
Figure 25: Evaluation performance of Point_Pillar
(a) Point_1Pillar
(b) Point_4Pillar
(c) Point_8Pillar
Figure 27: Training performance of Point_Pillar
(a) Point_1Pillar
(b) Point_4Pillar
(c) Point_8Pillar
Figure 29: ISSA performance of Point_Pillar
(a) Swimmer_1Hazard
(b) Ant_1Hazard
(c) Walker_1Hazard
Figure 31: Evaluation performance of link robots
(a) Swimmer_1Hazard
(b) Ant_1Hazard
(c) Walker_1Hazard
Figure 33: Training performance of link robots
(a) Swimmer_1Hazard
(b) Ant_1Hazard
(c) Walker_1Hazard
Figure 35: ISSA performance of link robots

### v4

@@ D.1 Preliminaries
Proof D.2.
Multiply both sides of (35) by $(I-\gamma P_{\pi})$ and take the inner product with the vector $f\in\mathbb{R}^{|\hat{\mathcal{S}}|}$.
Combining Lemma D.1 with (34), we obtain the following, for any function $f$ and any policy $\pi$:
@@
After bringing (43), $\left\|\dagger{\delta}_{f}^{\pi^{\prime}}\right\|_{\infty}$ into (42), then substract $\underset{\begin{subarray}{c}\hat{s}\sim\dot{d}^{\pi}\\ a\sim\pi\\ \hat{s}^{\prime}\sim P\end{subarray}}{E}[\delta_{f}(\hat{s},a,\hat{s}^{\prime})]$, the bounds are obtained. The lower bound leads to (39), and the upper bound leads to (40).
Lemma D.5.
@@
Proof D.8.
Begin with the bounds from lemma D.3 and bound the divergence $D_{TV}(\dot{d}^{\pi^{\prime}}||\dot{d}^{\pi})$ by lemma D.5.
@@ D.2 Safety invariance in expectation between S-3PO policies
Corollary D.9 (SCPO Update Constraint Satisfaction).
@@
Proof D.10.
The choice of $f=\hat{V}_{D}^{\pi},g=D$ in lemma D.7 leads to following inequality:
@@
where $\mathcal{\hat{J}}_{D}(\pi)=\mathbb{E}_{\tau\thicksim\pi}\Bigg{[}\sum_{t=0}^{H}\gamma^{t}D\big{(}\hat{s}_{t},a_{t},\hat{s}_{t+1}\big{)}\Bigg{]}$, need to distinguish from $\mathcal{J}_{D}(\pi)$. And $\hat{V}_{D}^{\pi},\hat{A}_{D}^{\pi}$ are also the discounted version of $V_{D}^{\pi}$ and $A_{D}^{\pi}$. Note that according to Lemma D.7 one can only get this the inequality holds when $\gamma\in(0,1)$.
@@
where $\bar{d}^{\pi}=\sum_{t=0}^{H}P(\hat{s}_{t}=\hat{s}|\pi)$. Thus, following the inequality (32), the Corollary D.9 is proofed.
D.3 Proof of Theorem 5.2
If $s_{t}\in\mathcal{\bar{S}}\in\mathcal{S}_{S}^{D}$, then $\mathcal{J}_{D}(\pi_{t})\leq 0$ and $\phi(s_{t})\leq 0$. According to Corollary D.9, we know that if policy $\pi_{t}$ is updated by solving (32), then we have $\mathcal{J}_{D}(\pi_{t+1})\leq 0$ which means:
@@
Appendix E Proof of Theorem 5.3
@@ E.2 Trust Region Update Performance
Proof E.4.
The choice of $f=V_{\pi},g=R$ in lemma D.7 leads to following inequality:
@@
At this point, the lemma E.3 is proved.
E.3 Proof of Theorem 5.3
Proof E.5.
If ISSA will constantly be triggered, the safeguard should be treated as a component of the environment, which indicates the environment is non-stationary for policy $\pi$ after safeguard is disabled. According to Algorithm 1, we assume that ISSA will not be triggered after $k_{safe}$ updates, meaning the environment is stationary for policy $\pi$ with or without safeguard. Therefore, we can infer S-3PO worst performance degradation after $k_{safe}$ updates following the trust region results of finite-horizon MDPs. Utilizing Lemma E.3 and the relationship between the total variation divergence and the KL divergence, we have:
@@
Due to Lemma E.1, if two policies are related with Equation 56, they are related with the following optimization:
@@ F.1 Environment Settings
To scale our method to high dimensional link robots. We additionally adopt Walker shown in 13 as a bipedal robot ($\mathcal{A}\subseteq\mathbb{R}^{10}$) in our experiments.
@@ F.2 Policy Settings
F.3 Practical Implementation Details
Weighted loss for cost value targets
Figure 14: $V_{[D_{i}]\pi}(\hat{s})$ target of five sampled episodes.
Line search scheduling
F.4 Metrics Comparison
In Tables 5, 6 and 7, we report all the $9$ results of our test suites by three metrics:
@@ F.4 Metrics Comparison
Both the evaluation performance and training performance are reported based on the above metrics. Besides, we also report the ISSA triger times as ISSA performance of TRPO-ISSA and S-3PO. All of the metrics were obtained from the final epoch after convergence. Each metric was averaged over two random seed. The evaluation performance curves of all experiments are shown in LABEL:fig:exp-point-hazard-eval, LABEL:fig:exp-point-pillar-eval and LABEL:fig:exp-link-robot-eval, the training performance curves of all experiments are shown in LABEL:fig:exp-point-hazard-train, LABEL:fig:exp-point-pillar-train and LABEL:fig:exp-link-robot-train and the ISSA performance curves of all experiments are shown in LABEL:fig:exp-point-hazard-ISSA, LABEL:fig:exp-point-pillar-ISSA and LABEL:fig:exp-link-robot-ISSA